# Cross-Domain Few-Shot Learning for Hyperspectral Image Classification Based on Mixup Foundation Model

### v1

@@ V-C Implementation Details
SSRN [68] 91.21 93.94 55.30 95.21 99.32 45.06 77.42 73.58 95.44 $76.78\pm 2.01$ $80.08\pm 2.54$ $69.32\pm 2.52$
MCNN [67] 81.23 86.31 57.62 94.23 99.13 72.82 77.41 69.63 95.57 $77.23\pm 2.25$ $80.89\pm 2.74$ $70.25\pm 2.62$
SPRN [64] 82.63 86.85 58.88 96.13 99.03 75.97 78.64 67.61 96.81 $79.52\pm 1.74$ $82.29\pm 1.49$ $73.64\pm 2.01$
DCFSL [24] 80.37 85.49 63.74 94.27 99.51 76.30 79.42 59.94 98.67 $81.52\pm 2.72$ $81.97\pm 1.23$ $76.10\pm 2.71$
Gia-CFSL [65] 88.12 85.95 66.44 92.78 99.18 75.03 74.58 72.42 94.67 $83.60\pm 2.29$ $83.21\pm 1.86$ $78.67\pm 2.73$
FDFSL [38] 86.07 83.09 65.16 89.29 99.40 75.11 88.66 85.19 78.34 $83.31\pm 1.89$ $85.19\pm 2.23$ $78.34\pm 2.16$
ADAFSL [57] 87.81 86.92 59.96 88.58 99.36 78.47 86.20 82.26 99.92 $85.12\pm 1.99$ $85.50\pm 1.78$ $80.59\pm 2.41$
RPCL-FSL [28] 87.57 87.53 67.62 94.44 99.02 59.45 85.88 86.89 99.36 $86.44\pm 1.97$ $85.86\pm 2.19$ $82.05\pm 1.73$
CASCL [25] 89.50 86.44 75.64 93.34 99.48 85.21 91.18 86.43 99.52 $89.75\pm 1.34$ $91.32\pm 1.86$ $85.52\pm 1.96$
GLGAT-CFSL [7] 95.85 96.95 85.09 91.92 98.8 92.57 92.28 90.95 98.98 88.17 91.71 88.35
CDLA [60] 88.48 84.38 80.92 90.89 98.84 86.37 97.77 88.9 97.27 $87.09\pm 4.35$ $90.42\pm 1.87$ $83.40\pm 5.13$
DA-CFSL [39] 92.95 87.13 83.71 88.78 99.6 84.44 97.65 90.47 96.44 $88.88\pm 2.70$ $91.24\pm 2.57$ $85.57\pm 3.35$
CFSL-KT [17] 98.79 86.36 86.71 93.95 99.25 96.92 99.92 73.78 94.42 $89.92\pm 2.50$ $92.22\pm 1.93$ $87.03\pm 3.11$
MGPDO [42] 92.09 92.50 89.85 96.14 99.46 98.71 98.91 90.75 99.61 $91.08\pm 1.07$ $94.66\pm 0.91$ $88.50\pm 1.37$
@@
TABLE VIII: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Salinas dataset (Five labeled samples per class)
@@
SSRN [68] 98.53 99.67 96.87 99.31 95.79 99.66 99.90 66.17 99.79 90.73 97.34 99.88 99.01 98.13 81.47 94.43 $89.13\pm 2.59$ $94.79\pm 0.55$ $87.95\pm 2.80$
MCNN [67] 98.41 99.76 91.96 99.25 92.70 99.52 98.58 74.57 99.59 83.42 96.61 98.93 98.30 98.24 76.38 92.22 $89.34\pm 2.19$ $94.04\pm 1.14$ $88.17\pm 2.41$
SPRN [64] 99.13 99.69 97.01 99.36 94.58 99.47 99.53 78.74 99.13 84.98 96.99 99.13 98.34 98.36 80.54 91.03 $90.05\pm 1.53$ $94.11\pm 0.96$ $88.63\pm 1.91$
DCFSL [24] 99.45 98.82 97.31 99.26 91.52 99.84 99.69 79.48 99.32 86.78 97.01 99.17 97.85 98.43 75.73 92.75 $90.51\pm 0.99$ $94.52\pm 0.66$ $89.45\pm 1.09$
Gia-CFSL [65] 98.09 99.41 95.46 98.52 86.55 99.25 99.28 81.31 97.76 83.73 93.56 99.19 94.27 98.46 74.51 90.32 $89.75\pm 0.80$ $93.10\pm 1.19$ $88.60\pm 0.90$
FDFSL [38] 98.96 98.78 80.10 99.32 90.47 98.41 99.76 78.85 97.98 85.40 95.51 99.64 98.29 99.00 82.53 92.04 $90.51\pm 1.17$ $93.94\pm 1.26$ $89.45\pm 1.30$
CF-Trans [70] 99.08 99.81 95.38 99.44 92.32 99.13 99.47 80.06 98.97 87.31 98.53 99.76 98.78 98.82 79.18 91,27 $91.03\pm 1.25$ $94.74\pm 0.69$ $90.03\pm 1.38$
ADAFSL [57] 99.19 99.68 97.10 99.57 95.74 99.26 99.30 77.67 99.66 84.64 99.08 99.92 99.09 99.44 84.40 92.62 $91.43\pm 0.47$ $95.28\pm 0.58$ $90.49\pm 0.52$
RPCL-FSL [28] 98.95 99.09 96.09 99.09 90.83 98.76 99.17 82.24 99.94 90.00 99.13 98.36 97.09 97.50 84.59 93.14 $92.33\pm 0.78$ $95.22\pm 0.71$ $91.47\pm 0.83$
CASCL [25] 99.45 99.62 99.84 99.49 95.94 99.82 99.91 84.52 99.71 92.69 98.30 99.81 99.66 98.24 83.91 94.23 $93.64\pm 0.63$ $95.89\pm 0.42$ $93.07\pm 0.69$
CDLA [60] 99.87 99.77 92.37 99.67 96.0 99.61 98.66 83.55 99.98 89.95 99.70 99.71 99.48 99.43 82.05 91.47 $92.62\pm 1.38$ $95.72\pm 0.95$ $91.79\pm 1.53$
CFSL-KT [17] 97.51 99.38 99.85 99.64 98.10 99.44 100.0 84.67 99.98 92.80 99.81 100.0 99.67 99.72 93.31 100.0 $95.13\pm 0.93$ $97.69\pm 0.52$ $94.60\pm 1.03$
MGPDO [42] 100 100 99.81 99.81 97.27 99.75 99.63 86.81 100 93.39 99.92 99.75 99.58 99.64 85.98 94.23 $94.19\pm 0.94$ $96.99\pm 0.40$ $93.54\pm 1.05$
@@
TABLE IX: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Houston dataset (Five labeled samples per class)
@@
SSRN [68] 73.00 83.98 92.83 80.95 94.44 77.91 67.28 25.47 81.09 52.55 49.67 25.03 81.47 99.27 98.76 $68.23\pm 4.21$ $72.25\pm 4.16$ $65.44\pm 4.75$
Gia-CFSL [65] 88.80 82.19 94.78 89.90 94.32 73.72 63.33 42.66 59.17 63.98 62.50 72.52 88.47 93.31 92.41 $5.05\pm 1.62 77$$7.47\pm 1.69$ $73.05\pm 1.74$
CNN-IT [16] 82.41 78.80 98.86 77.14 93.49 90.81 74.04 53.83 67.76 93.46 84.76 68.84 53.99 99.43 99.79 $79.58\pm 2.85$ $81.16\pm 2.62$ $77.92\pm 3.08$
HFSL [50] 89.61 79.43 99.65 90.72 91.09 90.97 89.93 44.02 68.67 68.66 68.76 62.84 68.38 96.47 96.77 $77.94\pm 2.38$ $80.20\pm 2.27$ $76.16\pm 2.56$
CMFSL [54] 89.74 87.47 95.78 92.28 97.38 79.93 73.76 48.04 70.12 69.35 66.73 70.94 95.60 96.00 97.72 $79.63\pm 2.60$ $82.06\pm 1.40$ $77.99\pm 2.21$
DCFSL [24] 89.13 83.25 97.12 92.11 97.09 80.72 70.13 41.92 64.15 63.23 62.15 74.42 88.04 91.58 97.66 $77.00\pm 1.71$ $79.51\pm 1.62$ $75.14\pm 1.86$
GMTN [49] 96.14 82.87 99.42 86.12 92.40 84.69 80.60 45.84 73.13 55.07 62.28 60.67 86.85 96.22 99.39 77.20 80.11 75.38
CMTL [5] 97.43 63.01 100.0 93.46 99.19 85.63 70.23 52.22 63.03 80.03 81.06 93.08 81.68 98.34 89.31 81.38 83.18 79.88
CF-Trans [70] 98.66 84.07 96.05 91.71 94.77 88.02 84.64 40.68 79.95 67.29 72.52 78.77 96.12 90.23 99.34 $82.02\pm 0.53$ $84.19\pm 0.78$ $80.57\pm 0.58$
CFSL-KT [17] 85.12 84.96 99.86 83.31 98.63 94.06 81.73 56.29 73.52 91.79 86.99 66.10 71.49 100.0 99.70 $82.63\pm 1.68$ $84.35\pm 1.70$ $81.23\pm 1.80$
@@
V-D Numerical Results
This subsection discusses our numerical results encompassing comparisons with recently published works. table VI reports numerical results of consolidated algorithms in the IP dataset as the target domain. The efficacy of our algorithm, MIFOMO, is observed where it outperforms other algorithms with a significant gap across all evaluation metrics, i.e.,  $15\%$ margin to MGPDO in the second place in the realm of overall accuracy (OA). This finding clearly demonstrates the efficacy of our mixup domain adaptation strategy combined with a foundational model. table VII shows numerical results of the consolidated algorithms in the PU dataset. It is seen that MIFOMO beats MGDPO in the second place with $6\%$ margin. table VIII exhibits numerical results of the consolidated algorithms, confirming the advantage of our algorithm as the best-performing method $1.5\%$ gap while table IX presents numerical results of the consolidated algorithms in the Houston dataset showing that MIFOMO outperforms CFSL-KT as the second-best method with a notable margin, $13\%$. This finding clearly substantiates our claim that MIFOMO produces state-of-the-art results over prior arts. MIFOMO features a hyper-spectral foundation model with a novel parameter fine-tuning approach, the coalescent projection. It is trained using the mixup domain adaptation method, which distinguishes itself from existing approaches, mostly using the adversarial-based domain adaptation method. That is, no domain discriminator is required in our framework. The mixup domain adaptation approach is underpinned by a two-phase training strategy, i.e., source and intermediate domains, where the intermediate domain functions as a bridge between the source domain and the target domain, enabling seamless knowledge transfer.
The classification maps of MIFOMO across four datasets are respectively displayed in Fig. 5, 6, 7, 8 and compared to the ground truth. It is seen that our algorithm is capable of delivering decent classification maps in respect to the ground truth. That is, the classification errors are low meaning that the predictions are close to the ground truth.
TABLE X: Ablation studies on the Indian Pines dataset.
@@ V-D Numerical Results
V-E Ablation Studies
Table X reports numerical results of our ablation study in the Indian Pines dataset, where different configurations of MIFOMO are tested and compared to demonstrate the influence of each learning module of MIFOMO. We commence our ablation study with the label smoothing module, refining the quality of pseudo-labels based on the consistency of neighboring samples. Obviously, the label smoothing approach plays a pivotal role in MIFOMO, where the absence of this module results in a notable performance drop, $17\%$. Note that the pseudo-labels are subsequently applied to create the intermediate domain. Unreliable pseudo labels lead to dramatic performance degradation. The intermediate domain training phase also demonstrates a key contribution to the success of MIFOMO, where its removal is attributed to performance deterioration by about $3\%$. This configuration implies that the model loses the target domain information. The parameter efficient fine-tuning (PEFT) technique via the coalescent projection strategy contributes to $1\%$ performance gain in the IP dataset. This configuration ignores the use of the PEFT technique, where the classification head is the only learnable module during the few-shot learning phase. Last but not least, the use of mixup for the data augmentation method and the establishment of the intermediate domain is linked to $2\%$ performance improvement. The mixup technique allows creations of reliable synthetic samples via linear interpolations of two distinct samples. Note that our mixup strategy is not only limited to the intra-class samples but also the inter-class samples. In other words, the model is trained to recognize new concepts or classes. The use of a foundation model alone doesn’t suffice for cross-domain few-shot learning in hyperspectral image classification. Although the foundation model already possesses generalizable features, certain techniques are required to achieve its best potential in downstream tasks.
V-F T-SNE Analysis

### v2

@@ 5.3 Implementation Details
SSRN Zhong2018SpectralSpatialRN 91.21 93.94 55.30 95.21 99.32 45.06 77.42 73.58 95.44 $76.78\pm 2.01$ $80.08\pm 2.54$ $69.32\pm 2.52$
MCNN zheng2020hyperspectral 81.23 86.31 57.62 94.23 99.13 72.82 77.41 69.63 95.57 $77.23\pm 2.25$ $80.89\pm 2.74$ $70.25\pm 2.62$
SPRN zhang2021spectral 82.63 86.85 58.88 96.13 99.03 75.97 78.64 67.61 96.81 $79.52\pm 1.74$ $82.29\pm 1.49$ $73.64\pm 2.01$
DCFSL Li2021DeepCF 80.37 85.49 63.74 94.27 99.51 76.30 79.42 59.94 98.67 $81.52\pm 2.72$ $81.97\pm 1.23$ $76.10\pm 2.71$
Gia-CFSL Zhang2022GraphIA 88.12 85.95 66.44 92.78 99.18 75.03 74.58 72.42 94.67 $83.60\pm 2.29$ $83.21\pm 1.86$ $78.67\pm 2.73$
FDFSL Qin2024CrossDomainFL 86.07 83.09 65.16 89.29 99.40 75.11 88.66 85.19 78.34 $83.31\pm 1.89$ $85.19\pm 2.23$ $78.34\pm 2.16$
ADAFSL Ye2023AdaptiveDF 87.81 86.92 59.96 88.58 99.36 78.47 86.20 82.26 99.92 $85.12\pm 1.99$ $85.50\pm 1.78$ $80.59\pm 2.41$
RPCL-FSL liu2023category 87.57 87.53 67.62 94.44 99.02 59.45 85.88 86.89 99.36 $86.44\pm 1.97$ $85.86\pm 2.19$ $82.05\pm 1.73$
CASCL Li2024CrossDomainFH 89.50 86.44 75.64 93.34 99.48 85.21 91.18 86.43 99.52 $89.75\pm 1.34$ $91.32\pm 1.86$ $85.52\pm 1.96$
GLGAT-CFSL Ding2024GLGATCFSLGG 95.85 96.95 85.09 91.92 98.8 92.57 92.28 90.95 98.98 88.17 91.71 88.35
CDLA Yu2025FewshotLF 88.48 84.38 80.92 90.89 98.84 86.37 97.77 88.9 97.27 $87.09\pm 4.35$ $90.42\pm 1.87$ $83.40\pm 5.13$
DA-CFSL Qin2025DualalignmentCF 92.95 87.13 83.71 88.78 99.6 84.44 97.65 90.47 96.44 $88.88\pm 2.70$ $91.24\pm 2.57$ $85.57\pm 3.35$
CFSL-KT Huang2023HyperspectralIC 98.79 86.36 86.71 93.95 99.25 96.92 99.92 73.78 94.42 $89.92\pm 2.50$ $92.22\pm 1.93$ $87.03\pm 3.11$
MGPDO Shi2025MultiGaussianPM 92.09 92.50 89.85 96.14 99.46 98.71 98.91 90.75 99.61 $91.08\pm 1.07$ $94.66\pm 0.91$ $88.50\pm 1.37$
@@
Table 8: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Salinas dataset (Five labeled samples per class)
@@
SSRN Zhong2018SpectralSpatialRN 98.53 99.67 96.87 99.31 95.79 99.66 99.90 66.17 99.79 90.73 97.34 99.88 99.01 98.13 81.47 94.43 $89.13\pm 2.59$ $94.79\pm 0.55$ $87.95\pm 2.80$
MCNN zheng2020hyperspectral 98.41 99.76 91.96 99.25 92.70 99.52 98.58 74.57 99.59 83.42 96.61 98.93 98.30 98.24 76.38 92.22 $89.34\pm 2.19$ $94.04\pm 1.14$ $88.17\pm 2.41$
SPRN zhang2021spectral 99.13 99.69 97.01 99.36 94.58 99.47 99.53 78.74 99.13 84.98 96.99 99.13 98.34 98.36 80.54 91.03 $90.05\pm 1.53$ $94.11\pm 0.96$ $88.63\pm 1.91$
DCFSL Li2021DeepCF 99.45 98.82 97.31 99.26 91.52 99.84 99.69 79.48 99.32 86.78 97.01 99.17 97.85 98.43 75.73 92.75 $90.51\pm 0.99$ $94.52\pm 0.66$ $89.45\pm 1.09$
Gia-CFSL Zhang2022GraphIA 98.09 99.41 95.46 98.52 86.55 99.25 99.28 81.31 97.76 83.73 93.56 99.19 94.27 98.46 74.51 90.32 $89.75\pm 0.80$ $93.10\pm 1.19$ $88.60\pm 0.90$
FDFSL Qin2024CrossDomainFL 98.96 98.78 80.10 99.32 90.47 98.41 99.76 78.85 97.98 85.40 95.51 99.64 98.29 99.00 82.53 92.04 $90.51\pm 1.17$ $93.94\pm 1.26$ $89.45\pm 1.30$
CF-Trans Zhu2025FromIT 99.08 99.81 95.38 99.44 92.32 99.13 99.47 80.06 98.97 87.31 98.53 99.76 98.78 98.82 79.18 91,27 $91.03\pm 1.25$ $94.74\pm 0.69$ $90.03\pm 1.38$
ADAFSL Ye2023AdaptiveDF 99.19 99.68 97.10 99.57 95.74 99.26 99.30 77.67 99.66 84.64 99.08 99.92 99.09 99.44 84.40 92.62 $91.43\pm 0.47$ $95.28\pm 0.58$ $90.49\pm 0.52$
RPCL-FSL liu2023category 98.95 99.09 96.09 99.09 90.83 98.76 99.17 82.24 99.94 90.00 99.13 98.36 97.09 97.50 84.59 93.14 $92.33\pm 0.78$ $95.22\pm 0.71$ $91.47\pm 0.83$
CASCL Li2024CrossDomainFH 99.45 99.62 99.84 99.49 95.94 99.82 99.91 84.52 99.71 92.69 98.30 99.81 99.66 98.24 83.91 94.23 $93.64\pm 0.63$ $95.89\pm 0.42$ $93.07\pm 0.69$
CDLA Yu2025FewshotLF 99.87 99.77 92.37 99.67 96.0 99.61 98.66 83.55 99.98 89.95 99.70 99.71 99.48 99.43 82.05 91.47 $92.62\pm 1.38$ $95.72\pm 0.95$ $91.79\pm 1.53$
CFSL-KT Huang2023HyperspectralIC 97.51 99.38 99.85 99.64 98.10 99.44 100.0 84.67 99.98 92.80 99.81 100.0 99.67 99.72 93.31 100.0 $95.13\pm 0.93$ $97.69\pm 0.52$ $94.60\pm 1.03$
MGPDO Shi2025MultiGaussianPM 100 100 99.81 99.81 97.27 99.75 99.63 86.81 100 93.39 99.92 99.75 99.58 99.64 85.98 94.23 $94.19\pm 0.94$ $96.99\pm 0.40$ $93.54\pm 1.05$
@@
Table 9: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Houston dataset (Five labeled samples per class)
@@
SSRN Zhong2018SpectralSpatialRN 73.00 83.98 92.83 80.95 94.44 77.91 67.28 25.47 81.09 52.55 49.67 25.03 81.47 99.27 98.76 $68.23\pm 4.21$ $72.25\pm 4.16$ $65.44\pm 4.75$
Gia-CFSL Zhang2022GraphIA 88.80 82.19 94.78 89.90 94.32 73.72 63.33 42.66 59.17 63.98 62.50 72.52 88.47 93.31 92.41 $5.05\pm 1.62 77$$7.47\pm 1.69$ $73.05\pm 1.74$
CNN-IT Huang2020HyperspectralIC 82.41 78.80 98.86 77.14 93.49 90.81 74.04 53.83 67.76 93.46 84.76 68.84 53.99 99.43 99.79 $79.58\pm 2.85$ $81.16\pm 2.62$ $77.92\pm 3.08$
HFSL Wang2022HeterogeneousFL 89.61 79.43 99.65 90.72 91.09 90.97 89.93 44.02 68.67 68.66 68.76 62.84 68.38 96.47 96.77 $77.94\pm 2.38$ $80.20\pm 2.27$ $76.16\pm 2.56$
CMFSL Xi2022FewShotLW 89.74 87.47 95.78 92.28 97.38 79.93 73.76 48.04 70.12 69.35 66.73 70.94 95.60 96.00 97.72 $79.63\pm 2.60$ $82.06\pm 1.40$ $77.99\pm 2.21$
DCFSL Li2021DeepCF 89.13 83.25 97.12 92.11 97.09 80.72 70.13 41.92 64.15 63.23 62.15 74.42 88.04 91.58 97.66 $77.00\pm 1.71$ $79.51\pm 1.62$ $75.14\pm 1.86$
GMTN Wang2023GraphMT 96.14 82.87 99.42 86.12 92.40 84.69 80.60 45.84 73.13 55.07 62.28 60.67 86.85 96.22 99.39 77.20 80.11 75.38
CMTL Cheng2023CausalML 97.43 63.01 100.0 93.46 99.19 85.63 70.23 52.22 63.03 80.03 81.06 93.08 81.68 98.34 89.31 81.38 83.18 79.88
CF-Trans Zhu2025FromIT 98.66 84.07 96.05 91.71 94.77 88.02 84.64 40.68 79.95 67.29 72.52 78.77 96.12 90.23 99.34 $82.02\pm 0.53$ $84.19\pm 0.78$ $80.57\pm 0.58$
CFSL-KT Huang2023HyperspectralIC 85.12 84.96 99.86 83.31 98.63 94.06 81.73 56.29 73.52 91.79 86.99 66.10 71.49 100.0 99.70 $82.63\pm 1.68$ $84.35\pm 1.70$ $81.23\pm 1.80$
@@
5.4 Numerical Results
This subsection discusses our numerical results encompassing comparisons with recently published works. table˜6 reports numerical results of consolidated algorithms in the IP dataset as the target domain. The efficacy of our algorithm, MIFOMO, is observed where it outperforms other algorithms with a significant gap across all evaluation metrics, i.e.,  $15\%$ margin to MGPDO in the second place in the realm of overall accuracy (OA). This finding clearly demonstrates the efficacy of our mixup domain adaptation strategy combined with a foundation model. table˜7 shows numerical results of the consolidated algorithms in the PU dataset. It is seen that MIFOMO beats MGDPO in the second place with $6\%$ margin. table˜8 exhibits numerical results of the consolidated algorithms, confirming the advantage of our algorithm as the best-performing method with $1.5\%$ gap while table˜9 presents numerical results of the consolidated algorithms in the Houston dataset showing that MIFOMO outperforms CFSL-KT as the second-best method with a notable margin, $13\%$. This finding clearly substantiates our claim that MIFOMO produces state-of-the-art results over prior arts. MIFOMO features a hyper-spectral foundation model with a novel parameter fine-tuning approach, the coalescent projection. It is trained using the mixup domain adaptation method, which distinguishes itself from existing approaches, mostly using the adversarial-based domain adaptation method. That is, no domain discriminator is required in our framework. The mixup domain adaptation approach is underpinned by a two-phase training strategy, i.e., source and intermediate domains, where the intermediate domain functions as a bridge between the source domain and the target domain, enabling seamless knowledge transfer.
The classification maps of MIFOMO across four datasets are respectively displayed in Fig. 5, 6, 7, 8, and compared to the ground truth. It is seen that our algorithm is capable of delivering decent classification maps with respect to the ground truth. That is, the classification errors are low, meaning that the predictions are close to the ground truth.
Table 10: Ablation studies on the Indian Pines dataset.
@@ 5.4 Numerical Results
5.5 Ablation Studies
Table 10 reports numerical results of our ablation study in the Indian Pines dataset, where different configurations of MIFOMO are tested and compared to demonstrate the influence of each learning module of MIFOMO. We commence our ablation study with the label smoothing module, refining the quality of pseudo-labels based on the consistency of neighboring samples. Obviously, the label smoothing approach plays a pivotal role in MIFOMO, where the absence of this module results in a notable performance drop, $17\%$. Note that the pseudo-labels are subsequently applied to create the intermediate domain. Unreliable pseudo labels lead to dramatic performance degradation. The intermediate domain training phase also demonstrates a key contribution to the success of MIFOMO, where its removal is attributed to performance deterioration by about $3\%$. This configuration implies that the model loses the target domain information. The parameter efficient fine-tuning (PEFT) technique via the coalescent projection strategy contributes to $1\%$ performance gain in the IP dataset. This configuration ignores the use of the PEFT technique, where the classification head is the only learnable module during the few-shot learning phase. Last but not least, the use of mixup for the data augmentation method and the establishment of the intermediate domain is linked to $2\%$ performance improvement. The mixup technique allows creations of reliable synthetic samples via linear interpolations of two distinct samples. Note that our mixup strategy is not only limited to the intra-class samples but also the inter-class samples. In other words, the model is trained to recognize new concepts or classes. The use of a foundation model alone doesn’t suffice for cross-domain few-shot learning in hyperspectral image classification. Although the foundation model already possesses generalizable features, certain techniques are required to achieve its best potential in downstream tasks.
5.6 T-SNE Analysis